# Conceptualization of this study, Methodology, Software

### v1

@@ 3 Experimental Setup
In this section, we describe the steps involved in the creation of the multimodal continual learning dataset as well as the standard evaluation criteria for the class incremental as well as domain incremental learning scenarios.
@@ 3.1 Continual Learning Scenario
According to the division of incremental batches and the availability of task identities, continual learning can be broadly classified into three primary types or scenarios Van de Ven et al.:
  • Task-Incremental Learning (TIL): It requires an algorithm to learn a set of distinct tasks progressively. Task identities/numbers are provided in both training and testing phase.
  • Class-Incremental Learning (CIL): The algorithm must learn to differentiate between a growing number of classes incrementally. Task identities are only provided during training phase.
  • Domain-Incremental Learning (DIL): The algorithm needs to learn the same tasks but in different domains. Task identities are not required.
In addition to these major scenarios, there are also other scenarios present in the literature like task-free continual learning (TFCL), online continual learning (OCL), blurred boundary continual learning (BBCL), continual pre-training (CPT), we refer to Wang et al. for a thorough study.
Figure 2: Commercial two fingers pneumatic gripper equipped with two force sensors and two flex sensors affixed together using stretchable nylon thread. The flex sensor measures the bending of the finger and the force sensor measures the magnitude of the force applied to the finger tip.
@@ 3.2 Multimodal Dataset
The multimodal dataset comprises two components: sensor signals from an actuated finger and non-stationary images of objects captured by a vision camera. The two different modalities are fused into a hybrid setup that incorporates both class incremental as well as domain incremental scenarios as illustrated in Figure 1. In this setup, progression from left to right represents class-incremental training, while progression from top to bottom represents domain-incremental training. Both domains encompass five tasks, with each task involving the learning of two new classes. The first domain focuses on sensor signal data for various objects encountered in each experience and follows a class-incremental learning approach. The second domain deals with visual data of the same objects and employs both domain-incremental and class-incremental learning methods. Further details on data collection for the different modalities are provided in the subsequent sections.
@@ 3.2.1 Sensor Signal
To collect the sensor signals for the different objects we utilize a soft pneumatic gripper Hao et al. equipped with four sensors which are affixed together using stretchable nylon thread as illustrated in Figure 10. We have employed the commercially available flex sensors Saggio et al., Mishra et al. to measure the bending of the finger and a force sensitive resistor (FSR) Hollinger and Wanderley to quantify the force applied to the fingertip as shown in Figure 2. The outputs from the flex and force sensors are directed to an Arduino Due board, which in turn is connected to a computer for further processing and analysis. We gather the sensor data for 10 distinct objects, for each object the gripper acquires 50 data points, considering various orientations of the object. During the data collection process, for each data point the gripper holds and releases the object repeatedly at different contact points, with each cycle lasting $\sim$ 39 seconds. Each data point is represented as a 600-dimensional vector, encapsulating the sensor signals recorded during that time interval. The dataset acquisition for every object takes about 37 minutes to complete on a computer equipped with an Intel Core i7 CPU @2.80GHz (GTX 1070 GPU) running Windows 10.
@@ 3.2.2 Image Data
3.3 Evaluation Metrics
The performance of the CL algorithm depends on its ability to adapt, retain, and generalize the knowledge over time. We evaluate the incremental accuracy (denoted as $a_{t}$) of the model after every task as well as the average accuracy over all the tasks for each domain denoted as $A_{T}$. The average accuracy over all the domains is referred to as $A_{D}$ in the paper. We define the expressions as follows:
$A_{T}=\frac{1}{T}\sum_{t=1}^{t=T}a_{t}\\ A_{D}=\frac{1}{D}\sum_{T=1}^{T=D}A_{T}$ (1)
@@ 3.3 Evaluation Metrics
To evaluate the SSL capabilities of the algorithm we perform two different experiments as proposed in the paper Smith et al., we use the following terminology to describe them: First, Random Images, where we have 20% of unlabeled data and 80% labeled data for each class. To collect the unsupervised data we randomly sample 20% (about 240) data from the training dataset for each class, remove the class label, shuffle, and store it in a separate folder. Second, Unseen Images, where we sample the images of an unseen object of the same parent class from the training dataset and repeat the above steps. A point to note is that in the first case, the model might have already seen all the objects for each class, however for the second case, the model has never seen the objects present in the unsupervised dataset and thus makes it a more challenging situation. In addition, we also perform an ablation study to show the effectiveness of the intra-layer feature representation as well as SSL on the model’s overall accuracy.
@@
Table 1: Pre-train model architecture for both the domains of the custom multimodal dataset, domain 2 represents the image data and domain 1 signifies the sensor data. The linear layers for both the models are removed once the pre-training is completed . Here, "classes" refers to the total number of output classes during the pre-training phase.
@@
Domain 2 Conv2D Relu 128 Batchnorm2D kernel:2, stride:2, padding:0
@@
4 Methodology
In this section, we discuss the modifications made to the FeCAM algorithm Goswami et al.. Originally, the algorithm models the feature distribution of classes using a multivariate normal distribution, $\mathcal{N}(\mu,\Sigma)$, and uses the Mahalanobis distance to perform classification. While the authors proposed several techniques to enhance and stabilize Mahalanobis-based distance classification, one significant limitation of FeCAM is its requirement to store batches in temporary memory and conduct training only once all data for a particular task is available. This makes it unsuitable for online learning. To address this issue and improve the algorithm’s capability for multimodal training, as well as to adapt to the semi-supervised nature of the dataset, we introduce several new features to the exFeCAM algorithm, which are detailed below.
4.1 Online Batchwise Training
Online training involves training a model incrementally in a batch-wise manner rather than on the entire dataset for a particular task at once. In this approach, each batch of data is passed through a feature extractor to obtain feature maps, which are then used to calculate the prototypes (mean and co-variance matrix) for each class in that batch. A dictionary is initialized with class labels as keys and their corresponding prototypes as values. There are two possible scenarios when updating this dictionary:
  1. 1.
    If a class is encountered for the first time, a new entry is added with the class label as the key and the prototypes as the value.
  2. 2.
This method avoids the need to store incoming data for each task, which is critical for continual learning (CL) objectives and is especially suitable for autonomous agents with limited cache memory. Learning prototypes in a batch-wise manner addresses the challenge of not having access to previous data points, making it a more practical solution for real-time applications.
@@ 4.1 Online Batchwise Training
Input: Data stream $(D_{0}^{0},D_{0}^{1},...,D_{1}^{0},D_{1}^{1},...,D_{n}^{0:t})$, where "$t$" refers to the number of tasks, "$n$" refers to the number of domains.
@@
Variables: $k$ : number of layers of feature extractor, $t$: total tasks, $n$: number of domains, $T$: total classes; $b$: batch size; $l,j$ : iterative variables
$(x_{i},y_{i})\in D_{train}^{Supervised}$, $(x_{i}^{{}^{\prime}})\in D_{train}^{Unsupervised}$ , $(x_{i}^{{}^{\prime\prime}},y_{i}^{{}^{\prime\prime}})\in D_{test}$
@@
Training Phase
for $E_{i}\in[E_{0},E_{1},...,E_{t}]$ do $\triangleright$ Different tasks/experiences
     Temporary Buffer $(\epsilon)\leftarrow\{\}$triangleright Temporary memory buffer for feature matching
     for $y_{i}\in[y_{0},y_{1},...,y_{T}]$ do $\triangleright$ Classes in each task
         for $j\leftarrow$ $b$ do
              $\psi_{ij}\leftarrow$ $[f_{e}(x_{i})_{k-(l-1)},f_{e}(x_{i})_{k-(l-2)},...,f_{e}(x_{i})_{k}]$triangleright Concatenate intra layer feature
              $\mu[i]\leftarrow\frac{1}{b}\sum\psi_{ij}$triangleright Mean matrix
              $\Sigma[i]\leftarrow$ Covariance($\psi_{ij}$) $\triangleright$ Covariance matrix
              $\Sigma[i]_{S}\leftarrow$ Shrinkage($\Sigma[i]$)
              $\Sigma[i]_{N}\leftarrow$ Normalization($\Sigma[i]_{S}$)
              if $\epsilon$ is not empty then
                  similarity $\leftarrow$ cosine similarity(features $\in\epsilon,x_{i}^{{}^{\prime}}$) $\triangleright$ Compute cosine similarity
                  if similarity $\geq$ threshold  then
                       Update $\mu[i]$, $\Sigma[i]_{N}$triangleright Update the mean and cov matrix for each class
                  end if
              end if
         end for
     end for
end for
Test Phase
for $x_{i}^{{}^{\prime\prime}}\in D_{test}$ do
     $\phi(f_{e}(x_{i}^{{}^{\prime\prime}});\mu,\Sigma)\leftarrow(f_{e}(x)-\mu_{i})^{T}$ $(\Sigma_{i})_{N}^{-1}$ $(f_{e}(x)-\mu_{i})$triangleright Calculate the Mahalanobis distance
     $y_{pred}\leftarrow$ argmin $\phi(f_{e}(x_{i}^{{}^{\prime\prime}});\mu,\Sigma)$ $\forall$ $y_{i}=0,1,...T$triangleright Assign the class with minimum Mahalanobis distance
end for
@@ 4.2 Intra Layer Feature Representation
Intra-layer feature representation (ILFR) leverages the power of intermediate layers that encodes the low and mid-level feature information present in the data extracted using a pre-trained model. Instead of relying solely on the final output layer of the feature extractor, the presence of multiple activation maps generalizes the learned knowledge and further improves the performance of the algorithm. We argue that enriching the last layer representations with hierarchical intra-layer features increases robustness to domain shifts and thus improves generalizability to downstream continual tasks. These extracted features can be the class embeddings of a transformer encoder Vaswani et al. or flattened feature maps of a ResNet encoder He et al..
There are two intuitive ways to promote intra-layer feature representations into the model: averaging the representations from the last "k" layers of the network or concatenating the representations from the last k layers. The first approach requires the output dimensions of all layers to be identical to facilitate summation. In our use case, we opt for concatenation, as the outputs from different layers have varying dimensions. Concatenation also preserves the rich, multi-scale features, unlike summation, which can result in information loss due to the merging of complementary features captured by different layers, potentially reducing the richness of the learned representations.
For the multimodal experiment, we pre-train the domain 1 feature extractor on gripper data of objects not present in the training dataset, while the domain 2 feature extractor is trained on the Core50 dataset. More details about the pre-trained model architecture are provided in Table 1. For the Core50 experiment, we use the ResNet18 architecture as the feature extractor, pre-trained on the ImageNet1k dataset Russakovsky et al..
@@ 4.3 Semi-Supervised Learning
Semi-supervised learning (SSL) denotes the capability of an algorithm to learn from supervised data and then improve the performance of the model by training it on the unsupervised part. It is very useful in a situation where we have limited supervised data but we can collect huge amounts of unsupervised data from the environment. This is an active area of research given that large, labeled datasets are expensive, but most applications have access to plentiful, cheap unlabeled data. In this paper, we use the concept of pseudo-labeling to organize the unsupervised data and provide them with dummy labels for further training. To perform pseudo labeling we randomly store the feature maps of each class for every task in a temporary memory buffer and then perform cosine similarity between the reference feature maps and the feature maps generated from the complete unsupervised dataset as defined in equation 2,
$similarity=\frac{x_{1}.x_{2}}{||x_{1}||_{2}||x_{2}||_{2}}$ (2)
where $x_{1}$ and $x_{2}$ refers to the reference feature map present in the temporary memory and the feature maps from the unsupervised dataset, ||.$||.||_{2}$ refers to the l2 norm of the two tensors. If the value of the cosine similarity is greater than a certain threshold, we then train the exFeCAM algorithm on the respective feature maps and its associated pseudo labels. These pseudo-labels are derived from the class labels used for feature matching. The prototype values for that class are subsequently updated using the averaging rule described in the previous section. To determine the threshold value, we employ the Optuna framework Akiba et al., which uses bayesian optimization to explore the hyperparameter space efficiently.
Figure 3: Modified FeCAM algorithm with intra layer feature representation, online batchwise learning, and semi-supervised learning capabilities for multimodal training. During training on multiple domains, the algorithm stores the mean and covariance matrices of the classes in a dictionary. In the testing phase, the algorithm uses these stored prototypes and employs the Mahalanobis distance to determine the prediction class. Supervised training is depicted with solid black lines, while unsupervised training is illustrated with dashed black lines.

### v2

@@ 3 Experimental Setup
In this section, we describe the steps involved in the creation of the multimodal continual learning dataset as well as the standard evaluation criteria for the class-incremental as well as domain-incremental learning scenarios.
@@ 3.1 Continual Learning Scenario
According to the division of incremental batches and the availability of task identities, continual learning can be broadly classified into three primary types or scenarios [46]:
  • Task-Incremental Learning (TIL): It requires an algorithm to learn a set of distinct tasks progressively, and it always knows which task it is working on.
  • Class-Incremental Learning (CIL): The algorithm must learn to differentiate between a growing number of classes incrementally.
  • Domain-Incremental Learning (DIL): The algorithm needs to learn the same task, but in different a domain or environment.
In addition to these major scenarios, there are also other scenarios present in the literature like task-free continual learning (TFCL), online continual learning (OCL), blurred boundary continual learning (BBCL), continual pre-training (CPT), we refer to [5] for a thorough study. In our experiment, we designed a hybrid scenario that integrates both class-incremental and domain-incremental learning paradigms.
Figure 2: Commercial two-finger pneumatic gripper equipped with two force sensors and two flex sensors affixed together using stretchable nylon thread. The flex sensor measures the bending of the finger and the force sensor measures the magnitude of the force applied to the fingertip.
@@ 3.2 Multimodal Dataset
The multimodal dataset comprises two components: sensor signal from an actuated finger and non-stationary images of objects captured by a vision camera. The two different modalities are structured into a CL setup that incorporates both class-incremental as well as domain-incremental scenarios as illustrated in Figure 1. In this setup, progression from left to right represents class-incremental training, while progression from top to bottom represents domain-incremental training. Both domains encompass five tasks, with each task involving the learning of two new classes. The first domain focuses on visual data for various objects encountered in each experience and follows a class-incremental learning approach. The second domain deals with sensor signal data of the same objects and employs both domain-incremental and class-incremental learning methods. Further details on data collection for the different modalities are provided in the subsequent sections.
In addition to our custom dataset, we also evaluate the effectiveness of our proposed method on a subset of the publicly available VGGSound dataset. We used this dataset due to it’s multimodal capabilities and it’s suitability for SSL settings. Specifically, we select five high-level parent classes: home, music, people, sports, and vehicles, each containing two specific subclasses. Every subclass includes 15 unique youtube videos for training, with each video lasting 240 seconds. Additionally, five separate videos per subclass are sampled in a similar fashion for testing. From these videos, we extract both the visual frames and the corresponding audio signals. To ensure accurate alignment between visual and audio modalities, we adopt the synchronization procedure outlined in [43]. Finally, the extracted visual frames are resized to 32x32 pixels to reduce computational complexity and accelerate the training process.
@@ 3.2.1 Sensor Signal
To collect sensor signals corresponding to different objects, we employ a soft pneumatic gripper [47] equipped with four sensors, which are secured in place using stretchable nylon thread. For capturing the bending motion of the fingers, we utilize commercially available flex sensors [48], [49], while fingertip force measurements are obtained using force-sensitive resistors (FSRs) [50], as illustrated in Figure 2. The outputs generated by both the flex and force sensors are routed to an Arduino Due board, which interfaces with a computer for subsequent signal processing and data analysis. We collect sensor readings across 10 different object classes, with each object contributing 50 individual data points that reflect variations in object orientation. During the data acquisition phase, the gripper holds and releases each object multiple times at different contact positions, with each cycle spanning approximately 39 seconds. Every data point is represented as a 600-dimensional vector that encapsulates the full set of sensor signals recorded within that duration. The complete data acquisition process for each object requires approximately 37 minutes and is carried out on a workstation equipped with an Intel Core i7 CPU @2.80GHz and a (GTX 1070 GPU).
@@ 3.2.2 Image Data
3.3 Evaluation Metrics and Benchmark Algorithms
The performance of the CL algorithm depends on it’s ability to adapt, retain, and generalize the knowledge over time. We evaluate the incremental accuracy (denoted as $a_{t}$) of the model after every task as well as the average accuracy over all the tasks for each domain denoted as $A_{T}$. The average accuracy over all the domains is referred to as $A_{D}$ in the paper. We define the expressions as follows:
$A_{T}=\frac{1}{T}\sum_{t=1}^{t=T}a_{t}\hskip 28.45274ptA_{D}=\frac{1}{D}\sum_{T=1}^{T=D}A_{T}$ (1)
@@ 3.3 Evaluation Metrics and Benchmark Algorithms
Additionally, to evaluate the SSL capability of our algorithm we perform two separate experiments as proposed in the paper [51], we use the following terminology to describe them: First, Random Images, where we have 30% of unlabeled data and 70% labeled data for each class. The unlabeled data is obtained by randomly sampling instances from the training dataset (ensuring that these images are removed from the labeled training set), removing their class labels, shuffling them, and storing them separately. Second, Unseen Images, where we sample the images of an unseen object of the same parent class from the training dataset and repeat the above steps. A point to note is that in the first case, the model might have already seen all the objects for each class however, for the second case, the model has never seen the objects present in the unsupervised dataset which makes it a more challenging situation.
To assess the effectiveness of the proposed exFeCAM algorithm, we compare it against five benchmark methods: naive-concat, naive-add, vanilla FeCAM, FeTrIL, and a joint training approach. The naive-concat method employs a non-CL strategy by concatenating feature maps from both modalities for each class, followed by incremental training on newly introduced objects. In contrast, naive-add applies the same incremental learning procedure but replaces concatenation with a weighted summation of the feature maps. Both naive-concat and naive-add utilize an identical 3-layer feedforward neural network architecture, but without any CL capability. Vanilla FeCAM corresponds to the original FeCAM architecture introduced by [8]. FeTrIL [52] also follows a replay free approach and stores the centroid representations of the past classes to produce their pseudo features by using geometric translation when training on new classes/domains. Lastly, the joint training approach involves offline training on the entire dataset, establishing an upper bound on accuracy.
Figure 3: exFeCAM algorithm with intra-layer feature representation, online batchwise learning, and SSL capabilities for multimodal training. During training on multiple domains, the algorithm stores the mean and covariance matrices of the classes in a dictionary. In the testing phase, the algorithm uses these stored prototypes and employs the Mahalanobis distance to determine the prediction class. Supervised training is depicted with solid black lines, while unsupervised training is illustrated with dashed black lines.
4 Methodology
In this section, we discuss the modifications made to the FeCAM algorithm [8]. Originally, the algorithm models the feature distribution of each class using a multivariate normal distribution, $\mathcal{N}(\mu,\Sigma)$, and employs a Mahalanobis distance based metric for classifying the respective samples. Although the authors proposed several techniques (co-variance matrix approximation, shrinkage and normalization) to enhance and stabilize the training process, one significant limitation of FeCAM is it’s requirement to store batches in temporary memory and conduct training only once all data for a particular task becomes available. This makes it unsuitable for online learning. To address this issue and improve the algorithm’s capability for multimodal training, as well as to adapt to the semi-supervised nature of the dataset, we introduce several new features to the exFeCAM algorithm, which are detailed below.
4.1 Online Batchwise Training
Online training refers to incrementally updating a model in a batch-wise fashion, rather than processing the entire dataset for a given task at once. In this approach, each batch of data is passed through a feature extractor to obtain feature maps, which are then used to calculate the prototypes (mean and co-variance matrix) for each class in that batch. A dictionary is initialized with class labels as keys and their corresponding prototypes as values. There are two possible scenarios when updating this dictionary:
  1. 1.
    If a class is encountered for the first time, a new entry is added with the class label as the key and the corresponding prototypes as the value.
  2. 2.
This method avoids the need to store incoming data for each task, which is critical for CL objectives and is especially suitable for autonomous agents with limited cache memory. Learning prototypes in a batch-wise manner addresses the challenge of not having access to previous data points, making it a more practical solution for real-time applications.
@@ 4.1 Online Batchwise Training
Table 1: Pre-train model architecture for both the domains of the custom multimodal dataset, domain 1 represents the image data and domain 2 signifies the sensor data. The linear layers for both the models are removed once the pre-training is completed . Here, ”classes” refers to the total number of output classes during the pre-training phase.
@@
Domain 1 Conv2D Relu 128 Batchnorm2D kernel:2, stride:2, padding:0
@@
Input: Data stream $\left(D_{0}^{0},D_{0}^{1},\dots,D_{1}^{0},D_{1}^{1},\dots,D^{0:t}_{n}\right)$
@@
Variables: $k$: number of layers in feature extractor, $t$: total tasks, $n$: number of domains, $T$: total classes, $b$: batch size, $l,j$: iterative variables
   $(x_{i},y_{i})\in D_{\text{train}}^{\text{Supervised}}$, $(x_{i}^{{}^{\prime}})\in D_{\text{train}}^{\text{Unsupervised}}$, $(x_{i}^{{}^{\prime\prime}},y_{i}^{{}^{\prime\prime}})\in D_{\text{test}}$
   $\mu,\Sigma\leftarrow\{\}$
Training Phase:
1:for each experience $E_{i}\in\{E_{0},E_{1},\dots,E_{t}\}$ do
2:  Temporary Buffer $(\epsilon)$ $\leftarrow\{\}$ {Temporary memory buffer for feature matching}
3:  for each class $y_{i}\in\{y_{0},y_{1},\dots,y_{T}\}$ do
4:   for $j=0$ to $b$ do
5:    $\psi_{ij}\leftarrow\large[{f_{e}}{(x_{i})}_{k-(l-1)},{f_{e}}{(x_{i})}_{k-(l-2)},\dots,f_{e}(x_{i})_{k}$ ] {Concatenate intra-layer features}
6:    $\mu[i]\leftarrow\frac{1}{b}\sum\psi_{ij}$ {Mean matrix}
7:    $\Sigma[i]\leftarrow\text{Covariance}(\psi_{ij})$ {Covariance matrix}
8:    $\Sigma[i]_{S}\leftarrow\text{Shrinkage}(\Sigma[i])$
9:    $\Sigma[i]_{N}\leftarrow\text{Normalization}(\Sigma[i])_{S}$
10:    if $\epsilon$ not empty then
@@
12:     if similarity $\geq$ threshold then
13:      Update $\mu[i]_{S},\Sigma[i]_{N}${Update the mean and cov matrix for each class}
14:     end if
15:    end if
16:   end for
17:  end for
18:end for
Test Phase:
1:for each $x_{i}^{{}^{\prime\prime}}\in D_{\text{test}}$ do
2:  $\phi(f_{e}(x_{i}^{{}^{\prime\prime}});\mu,\Sigma)\leftarrow(f_{e}(x)-\mu_{i})^{T}(\sum_{i})_{N}^{-1}(f_{e}(x)-\mu_{i})$ {Calculate the Mahalanobis distance}
3:  $y_{\text{pred}}\leftarrow\arg\min\phi(f_{e}(x_{i}^{{}^{\prime\prime}});\mu,\Sigma),\forall y_{i}\in\{0,1,\dots,T\}${Assign the class with minimum Mahalanobis distance}
4:end for
@@ 4.2 Intra Layer Feature Representation
Intra-layer feature representation (ILFR) leverages the power of intermediate layers that encodes the low and mid-level feature information present in the data extracted using a pre-trained model. Instead of relying solely on the final output layer of the feature extractor, the presence of multiple activation maps generalizes the learned knowledge and further improves the performance of the algorithm. We argue that enriching the last layer representations with hierarchical intra-layer features increases robustness to domain shifts and thus improves generalizability to downstream continual tasks. These extracted features can be the class embeddings of a transformer encoder [53] or the flattened feature maps of a ResNet encoder [54].
There are two intuitive ways to promote intra-layer feature representations into the model: averaging the representations from the last ”k” layers of the network or concatenating the representations from the last ”k” layers. The first approach requires the output dimensions of all layers to be identical to facilitate summation. In our use case, we opt for concatenation, as the outputs from different layers have varying dimensions. Concatenation also preserves the rich, multi-scale features, unlike summation, which may result in information loss due to the merging of complementary features captured by different layers, potentially reducing the richness of the learned representations.
For the multimodal experiment, we pre-train the domain 1 feature extractor on gripper data of objects not present in the training dataset, while the domain 2 feature extractor is pre-trained on the Core50 dataset [55]. For the VGGSound experiment, we use the ResNet18 architecture as the feature extractor, pre-trained on the ImageNet1k dataset [56]. More details about the pre-trained model architecture are provided in Table 1.
Figure 4: On the left: intra layer feature representation for the custom dataset and the VGGSound dataset, we compare the improvement in the accuracy of the FeCAM algorithm on addition of knowledge from the last as well the penultimate layers of the pre-trained feature extractors. On the right: the plot shows sensor data accuracy (in orange) with and without the addition of random gaussian noise, the solid line in black shows the improvement in the prediction accuracy for the different experiences after addition of the image domain.
@@ 4.3 Semi-Supervised Learning
We incorporate an SSL approach to enable our algorithm to leverage unlabeled data, thus enhancing it’s overall generalization and improving performance on both previously encountered and newly introduced objects. To integrate SSL capabilities into the exFeCAM algorithm, we adopt a pseudo-labeling strategy, which allows us to organize the unlabeled data by assigning them dummy labels for further training. The pseudo-labeling process begins by randomly storing few feature maps of each class for every task in a temporary memory buffer. We then compute the cosine similarity between these reference feature maps and the feature maps generated from incoming unlabeled data, using the formula defined in Equation 2.
[…]
Here, $x_{1}$ represents a reference feature map from the temporary memory buffer, while $x_{2}$ corresponds to the feature map extracted from the new unlabeled data. The notation $||.||_{2}$ indicates the L2 norm of the respective vectors. If the cosine similarity between the two feature maps exceeds a predefined threshold, the exFeCAM algorithm proceeds to train on the new feature map and it’s assigned pseudo-label. These pseudo-labels are derived from the class labels associated with the most similar reference feature used during feature matching. The corresponding prototype values for that class, specifically the mean and covariance matrices are then updated according to the averaging rule discussed in the previous section.
To select an appropriate threshold for pseudo-labeling, we utilize the Optuna framework [57], which employs bayesian optimization to efficiently explore the hyperparameter space and identify an optimal value. Once the pseudo-labeling process for a particular task is complete, the temporary memory buffer is discarded.
Table 2: Comparison of the accuracy of the exFeCAM algorithm and the benchmark methods across two different multimodal datasets. The joint method represents a scenario where the algorithm has access to the entire dataset at once.
Dataset Joint Naive-add Naive-concat FeTrIL Vanilla FeCAM ExFeCAM
Custom multimodal 94.65±1.29 43.43±1.59 45.16±1.68 62.10±1.86 65.30±0.98 70.50±1.79
VGGSound 92.48±2.31 29.78±1.76 30.94±1.81 38.21±0.89 39.96±1.16 45.96±1.98